# Mixed-wet percolation on a dual square lattice

### v1

@@ 4 Results and Discussion
Figure 5: (a) Plot of $W_{b}(p,L)$ versus $p$ for different $L$. All curves intercept at $p=p_{c}=0.4072$ as shown by a vertical dotted line. The dashed horizontal line represents $W_{b}(p,L)=1/2$. The intersection of $W_{b}(p,L)$ with $W_{b}(p,L)$ occurs at the local threshold $p_{c}(L)$ of individual $L$. The coloured bullets on the $x$-axis represent $p_{c}(L)$ of the respective $L$. (b) Plot of $p_{c}(L)$ versus $L^{-1/\nu}$ taking $\nu=4/3$, the correlation length exponent of ordinary site percolation . The straight line represents the regression line. The intercepts with the $y$-axis, denoted by a red cross, yield $p_{c}=0.4072\pm 0.0034$.
Figure 6: Plot of $W^{\prime}_{b}(p_{c},L)$ versus $L$ in double logarithmic scale is shown by circles. The straight line through the circles represents the regression line with slope $0.741\pm 0.002$. In the inset, $\log\xi$ is plotted against $\log|p-p_{c}|$ for $L=2048$ in the range $p<p_{c}$, shown by squares. The straight line through the squares is the regression line with slope $\nu=1.332\pm 0.001$.
4.1 Simulation
An extensive computer simulation of the above model has been performed on a square lattice and its dual lattice of different sizes $L\times L$, varying from $L=200$ to $2000$ in steps of $200$. Averaging of the geometrical quantities is made over $N=10^{5}$ samples. Periodic boundary conditions are applied in both directions. The burning algorithm [26] is used both in the horizontal and vertical directions to identify the spanning of the clusters. The Hoshen-Kopelman algorithm [27] is used to find the numbers of clusters $n_{b}$ with their sizes $b$ to determine the cluster size distributions and other geometrical properties of the cluster. In Fig.4, we present typical mixed wet percolation clusters generated on a square dual lattice of size $100\times 100$ (a) for site occupation probability $p=0.407$ on the primal lattice and (b) for bond occupation probability $p=0.5$ on the square lattice. Different colours represent clusters of certain sizes. The largest cluster is in red. It can be noticed that in Fig. 4(a) all clusters appear in the form of closed loops connected at certain corner points where four bonds meet. Moreover, in mixed wet percolation clusters (Fig.4(a)), the red bonds [28] are absent, whereas there are many such bonds in an ordinary random bond percolation cluster (Fig.4(b)). The bob-link model [29] of percolation is then not applicable to mixed wet percolation, contrary to ordinary bond percolation. Thus, the perimeter bond clusters of mixed wet percolation are very different from the ordinary bond percolation clusters [30] where one finds open dangling ends and red bonds. It is, therefore, important to characterize the critical properties of the mixed wet percolation clusters.
@@ 4.2 Wrapping probability: $p_{c}$ and $\nu$
$W_{b}(p_{c}(L),L)=\widetilde{W}_{b}\left[(p_{c}(L)-p_{c})L^{1/\nu}\right]=1/2$ (16)
@@
In Fig. 5 (b), the values of $p_{c}(L)$ are plotted against $L^{-1/\nu}$ taking $\nu=4/3$, the correlation length exponent of the ordinary percolation. A straight line is then fitted through the data points. Since the points follow a straight line, it indicates the value of $\nu$ is taken rightly. From the intercept of the fitted line with the $y$-axis, we obtain the percolation threshold of mixed wet percolation at $p_{c}=0.4072\pm 0.0033$ indicated by a cross on the $y$-axis. The error includes the least square fit error and the error due to uncertainty in $p_{c}(L)$. The $p_{c}$ obtained here is close to the next-nearest-neighbour (nnn) site percolation threshold ($\approx 0.4072$) on the square lattice [31]. Note that $0.407253791-p_{c}({\rm OSP})=$ 0.40725379 where $p_{c}({\rm OSP})=0.59274621$ is the ordinary site percolation (OSP) [32]. Thus, at $p_{c}$, along with mixed wet percolation in the dual lattice, nnn site percolation and breakdown of the infinite network of nearest-neighbour unoccupied sites occur on the primal lattice. It can be seen in Fig. 3(a) and (c) that a knot in mixed wet percolation in the dual lattice corresponds to an nnn bond on the primal lattice. The corresponding critical bond density $\rho_{c}$ on the dual lattice is $\rho_{c}=2p_{c}(1-p_{c})\approx 0.4828$, and it is below $0.5$, the random bond percolation threshold on the square lattice.
@@
$\displaystyle W^{\prime}_{b}(p_{c},L)=cL^{1/\nu}$ (18)
where $c=\widetilde{W}^{\prime}_{b}\left[0\right]$ is a constant, one can determine $1/\nu$ by measuring $W^{\prime}_{b}(p_{c},L)$ for different $L$. The slopes are measured right at the critical threshold $p_{c}$ by fitting a third-degree polynomial through ten data points around $p_{c}$ for each $L$. The values of $W^{\prime}_{b}(p_{c},L)$ are plotted against $L$ in a double logarithmic scale in Fig. 6. By linear regression through the data points, we found $1/\nu=0.741\pm 0.002$, which is very close to the ordinary site percolation value of $0.751/\nu=$ 0.75. The value of the exponent $\nu$ is further verified by estimating the correlation length $\xi$ (Eq. 11) as a function of $|p-p_{c}|$ on a lattice of size $L=2048$. In the inset, $\xi$ is plotted against $|p-p_{c}|$ in a double logarithmic scale. Through regression, we obtain $\nu=1.332\pm 0.001$. Thus, it shows that the correlation length exponent is the same as that of the ordinary site percolation, i.e.; $\nu=4/3$. Since the percolation criteria for the nnn percolation and perimeter bond percolation are the same, the wrapping probability $W(p,L)$ as a function $p$ for both problems will be the same. Hence, it is not surprising that the critical threshold $p_{c}$ and $1/\nu$ are the same for both problems.
Now, we investigate the general form of the wrapping probability in terms of $p$ on a large single lattice. In Fig. 7, $W_{b}(p)$ for $L=1000$ is plotted against $p$, varying $p$ from $0.3$ to $0.7$. It can be seen that the $W_{b}(p)$ is symmetric around $p=1/2$. The horizontal dotted line represents the value of $W_{b}(p_{c},L)$. This line intercepts the first dashed vertical line ($p_{c}=0.40725$) at the threshold point. It also intercepts $W_{b}(p)$ on the other side ($p>1/2$) at another threshold, which is indicated by another dashed line $0.592751-p_{c}=$ 0.59275. Two black crosses on the $x$-axis indicate these threshold values. Note that the second threshold, $0.592751-p_{c}=$ 0.59275, is the ordinary site percolation threshold $p_{c}({\rm OSP})$ [2]. Then the first $p_{c}$, is $p_{c}=1-p_{c}({\rm OSP})=0.40725$, the threshold of nnn site percolation.
Figure 7: Plot of $W_{b}(p,L)$ versus $p$ for the perimeter bond clusters on the dual lattice for a $1000\times 1000$ lattice. The black crosses on the $x$-axis indicate the thresholds $p_{c}$ and $1-p_{c}$.
@@
$\displaystyle W_{b}(p,\infty)$ $=H(p-p_{c})H\left[(1-p_{c})-p\right]$ (19)
@@
Figure 8: Plot of cluster size distribution of perimeter bond clusters $n_{b}(p_{c})$ against the size of the perimeter bond cluster $b$ and cluster size distribution of nnn sites clusters $n_{s}(p_{c})$ against nnn cluster size $s$ in a double logarithmic scale.
@@ 4.3 Cluster size distribution
$n_{b}(p_{c})\sim b^{-\tau}$ (20)
as $\mathcal{N}[0]$ is a constant. In Fig. 8, the cluster size distribution $n_{b}(p_{c})$ of the perimeter bond cluster is plotted against the cluster size $b$ for the system of size $L=2000$ in the double logarithmic scale shown by the red squares. For comparison, the size distribution of the nnn site percolation clusters $n_{s}(p_{c})\sim s^{\tau}$ is shown by blue triangles in Fig. 8. The slopes of the straight part of both the distributions are found to be $2.054\pm 0.002$. It is very close to the exponent $\tau=187/91$ as that of ordinary site percolation, shown by a black line as a guide to the eye. Note that the distributions $n_{b}(p_{c})$ and $n_{s}(p_{c})$ are two independent distributions. However, they only differ in the first bin and in the last few bins, though the total number of site and perimeter clusters are the same. In the first bin ($s\text{ or }b\in 1-10$), though there are all possible nnn site percolation clusters, the perimeter bond clusters appear in even numbers starting from $b=4$ only. The largest bin of $n_{b}(p_{c})$ is greater than that of $n_{s}(p_{c})$. This indicates that the perimeter of a site cluster can be more than the size of the site cluster.
Refer to caption Refer to caption
(a) (b)
Figure 9: (a) Plot of cluster size distribution of perimeter bond clusters $n_{b}(p)$ against perimeter cluster size $b$ in a double logarithmic scale for various values of $p$, indicated by different symbols, for a system size of $L=2048$. (b) Plot of $b_{\xi}$ and $s_{\xi}$ against $p-p_{c}$ in a double logarithmic scale. The value of critical exponent $\sigma=2.52\pm 0.03$ was determined.
Though the distribution is fairly power law for $p=p_{c}$, it deviates from the power law for $p<p_{c}$. The perimeter bond cluster size distributions for $p<p_{c}$ are plotted in Fig. 9(a) in double logarithmic scale. There is a strong deviation from the power law as $p$ is smaller than $p_{c}$. For a given $p$, if such a crossover occurs at a certain value of the perimeter bond cluster size $b_{\xi}(p)$, the cluster size distribution can written in terms of $b_{\xi}$ as in [33]
$n_{b}(p)=n_{b}(p_{c})F\left(\frac{b}{b_{\xi}}\right)$ (21)
where $n_{b}(p_{c})\sim b^{-\tau}$ and $b_{\xi}$ scales as
@@
$\sigma$ is an exponent. We have measured $b_{\xi}$ from the intersection of the curve $\log(n_{b}(p))$ with the line $\log(n_{b}(p_{c}))$. To find the intersection point, we fit a higher order polynomial through $\log{(n_{b}(p))}$ around the point of intersection and find the solution with the line of $\log(n_{b}(p_{c}))$. Red crosses mark the $\log b_{\xi}$ values on the $x$-axis. In Fig. 9(b), $b_{\xi}$ is plotted against $|p-p_{c}|$ in double logarithmic scale and found that $b_{\xi}$ follows a power law with an exponent $1/\sigma=2.52\pm 0.03$ which is very close to the percolation value $1/\sigma=91/36\approx 2.53$. The corresponding quantity in nnn site percolation $s_{\xi}$ is also plotted for comparison. It seems the value of $\sigma$ is the same as that of the ordinary percolation.
@@
is a normalized scaling function of the scaled variable $z=(p-p_{c})b^{\sigma}$. The data for different $p$ for a given $L$ should collapse onto a single curve if $n_{b}(p)/n_{b}(p_{c})$ is plotted against $z$. In Fig. 10, the ratio $n_{b}(p)/n_{b}(p_{c})$ is plotted against the scaled variable $z=(p-p_{c})b^{\sigma}$ for several values of $p$ for the system of $L=2000$ taking $\tau=187/91$ and $\sigma=36/91$ as that of the ordinary site percolation . It can be seen that data for several $p$ have collapsed to a single curve with the $\tau$ and $\sigma$ of the ordinary site percolation. It can be noted that at $p=p_{c}$, one finds $z=0$ and that corresponds to $\mathcal{N}[0]/\mathcal{N}[0]=1$ on the $y$-axis. In terms of the parameter $\rho$, $n_{b}(\rho)$ is indifferent from $n_{b}(p)$. However, the normalized plot of $n_{b}(\rho)/n_{b}(\rho_{c})$ against the scaled variable $z=(\rho-\rho_{c})b^{\sigma}$ will have the same height but squeezed width.
Figure 10: Plot of the collapse of ratio of cluster size distributions $n_{b}(p)/n_{b}(p_{c})$ against $(p-p_{c})b^{\sigma}$ for various values of $p$ at lattice size $L=2000$.

### v2

@@ 4 Results and Discussion
4.1 Simulation
An extensive computer simulation of the above model has been performed on a square lattice and its dual lattice of different sizes $L\times L$, varying from $L=200$ to $2000$ in steps of $200$. Averaging of the geometrical quantities is made over $N=10^{5}$ samples. Periodic boundary conditions are applied in both directions. The burning algorithm hans is used both in the horizontal and vertical directions to identify the spanning of the clusters. The Hoshen-Kopelman algorithm hk is used to find the number of clusters $n_{b}$ with their sizes $b$ to determine the cluster size distributions and other geometrical properties of the cluster. In Fig.4 (a), we present a typical mixed wet percolation cluster configuration generated on a square dual lattice of size $100\times 100$ for site occupation probability $p=0.407$ on the primal lattice. In Fig.4(b), a typical random bond cluster configuration is shown for bond occupation probability $p=0.5$ on the square (primal) lattice of size $100\times 100$ for comparison to the MWP cluster configuration in Fig.4(a). A particular colour is assigned to clusters of a certain range of cluster sizes. The ranges are fixed by logarithmic binning of sizes. The largest cluster is in red. It can be noticed that in Fig. 4(a) all clusters appear in the form of closed loops connected at certain corner points where four bonds meet. Moreover, in mixed wet percolation clusters (Fig.4(a)), the red bonds (singly connected bonds) de1986multiscaling are absent, whereas there are many such bonds in an ordinary random bond percolation cluster (Fig.4(b)). The bob-link model stanley1977cluster of percolation is then not applicable to mixed wet percolation, contrary to ordinary bond percolation. Thus, the perimeter bond clusters of mixed wet percolation are very different from the ordinary bond percolation clusters fan2021statistical , where one finds open dangling ends and red bonds. It is, therefore, important to characterize the critical properties of the mixed wet percolation clusters.
@@ 4.1 Simulation
Refer to caption Refer to caption
(c) (d)
Figure 5: (a) Plot of $W_{b}(p,L)$ versus $p$ for different $L$. All curves intercept at $p=p_{c}=0.4072$ as shown by a vertical dotted line. The dashed horizontal line represents $W_{b}(p,L)=1/2$. The intersection of $W_{b}(p,L)$ with $W_{b}(p,L)=1/2$ occurs at the local threshold $p_{c}(L)$ of individual $L$. The coloured bullets on the $x$-axis represent $p_{c}(L)$ of the respective $L$. (b) Plot of $p_{c}(L)$ versus $L^{-1/\nu}$ taking $\nu=4/3$, the correlation length exponent of ordinary site percolation. The straight line represents the regression line. The intercepts with the $y$-axis, denoted by a red cross, yield $p_{c}=0.4072\pm 0.0034$. (c) Plot of $W^{\prime}_{b}(p_{c},L)$ versus $L$ in a double logarithmic scale. The data points are represented by circles. The straight line through the circles represents the regression line with slope $0.741\pm 0.002$. In the inset, the logarithm of correlation length $\log\xi$ is plotted against $\log|p-p_{c}|$ for $L=2048$ in the range $p<p_{c}$, shown by squares. The straight line through the squares is the regression line with slope $\nu=1.332\pm 0.001$. (d) Plot of $W_{b}(p,L)$ versus $p$ for the perimeter bond clusters on the dual lattice for a $1000\times 1000$ lattice. The black crosses on the $x$-axis indicate the thresholds $p_{c}$ and $1-p_{c}$.
@@ 4.2 Wrapping probability: $p_{c}$ and $\nu$
$W_{b}(p_{c}(L),L)=\widetilde{W}_{b}\left[(p_{c}(L)-p_{c})L^{1/\nu}\right]=1/2\;,$ (16)
@@
In Fig. 5 (b), the values of $p_{c}(L)$ are plotted against $L^{-1/\nu}$ taking $\nu=4/3$, the correlation length exponent of the ordinary percolation. A straight line is then fitted through the data points. Since the points follow a straight line, it indicates the value of $\nu$ is taken rightly. From the intercept of the fitted line with the $y$-axis, we obtain the percolation threshold of mixed wet percolation at $p_{c}=0.4072\pm 0.0033$ indicated by a cross on the $y$-axis. The error includes the least square fit error and the error due to uncertainty in $p_{c}(L)$. The $p_{c}$ obtained here is close to the next-nearest-neighbour (nnn) site percolation threshold ($\approx 0.4072$) on the square lattice xun2021site . Here, by the term next-nearest-neighbour (nnn), we refer to the $8$ neighbours surrounding a site on a square lattice. These include the $4$ nearest-neighbour sites (left, right, top, and bottom) and the $4$ diagonal sites at a distance $\sqrt{2}$ times the lattice constant. Different studies have used varying terminologies for these neighbour classes. For example, in mg05 , they were denoted as SQ-1,2, whereas in nz00 , they were labeled as NN+NNN. Since this article does not discuss any additional nearest-neighbour sets, we will consistently use the term next-nearest-neighbour (nnn) to refer to the aforementioned eight neighbouring sites.
Note that $1-p_{c}({\rm OSP})=0.40725379$ where $p_{c}({\rm OSP})=0.59274621$ is the ordinary site percolation (OSP) nz00 . Thus, at $p_{c}$, along with mixed wet percolation in the dual lattice, nnn site percolation and breakdown of the infinite network of nearest-neighbour unoccupied sites occur on the primal lattice. It can be seen in Fig. 3(a) and (c) that a knot in mixed wet percolation in the dual lattice corresponds to an nnn bond on the primal lattice. The corresponding critical bond density $\rho_{c}$ on the dual lattice is $\rho_{c}=2p_{c}(1-p_{c})\approx 0.4828$, and it is below $0.5$, the random bond percolation threshold on the square lattice.
@@
$\displaystyle W^{\prime}_{b}(p_{c},L)=cL^{1/\nu}\;,$ (18)
where $c=\widetilde{W}^{\prime}_{b}\left[0\right]$ is a constant, one can determine $1/\nu$ by measuring $W^{\prime}_{b}(p_{c},L)$ for different $L$. The slopes are measured right at the critical threshold $p_{c}$ by fitting a third-degree polynomial through ten data points around $p_{c}$ for each $L$. The values of $W^{\prime}_{b}(p_{c},L)$ are plotted against $L$ in a double logarithmic scale in Fig. 5(c). By linear regression through the data points, we found $1/\nu=0.741\pm 0.002$, which is very close to the ordinary site percolation value of $1/\nu=0.75$. The value of the exponent $\nu$ is further verified by estimating the correlation length $\xi$ (Eq. 11) as a function of $|p-p_{c}|$ on a lattice of size $L=2048$. In the inset, $\xi$ is plotted against $|p-p_{c}|$ in a double logarithmic scale. Through regression, we obtain $\nu=1.332\pm 0.001$. Thus, it shows that the correlation length exponent is the same as that of the ordinary site percolation, i.e.; $\nu=4/3$. Since the percolation criteria for the nnn percolation and perimeter bond percolation are the same, the wrapping probability $W(p,L)$ as a function $p$ for both problems will be the same. Hence, it is not surprising that the critical threshold $p_{c}$ and $1/\nu$ are the same for both problems.
Now, we investigate the general form of the wrapping probability in terms of $p$ on a large single lattice. In Fig. 5(d), $W_{b}(p)$ for $L=1000$ is plotted against $p$, varying $p$ from $0.3$ to $0.7$. It can be seen that the $W_{b}(p)$ is symmetric around $p=1/2$. The horizontal dotted line represents the value of $W_{b}(p_{c},L)$. This line intercepts the first dashed vertical line ($p_{c}=0.40725$) at the threshold point. It also intercepts $W_{b}(p)$ on the other side ($p>1/2$) at another threshold, which is indicated by another dashed line $1-p_{c}=0.59275$. Two black crosses on the $x$-axis indicate these threshold values. Note that the second threshold, $1-p_{c}=0.59275$, is the ordinary site percolation threshold $p_{c}({\rm OSP})$ stauffer2018introduction . Then the first $p_{c}$, is $p_{c}=1-p_{c}({\rm OSP})=0.40725$, the threshold of nnn site percolation.
@@
$\displaystyle W_{b}(p,\infty)$ $=H(p-p_{c})H\left[(1-p_{c})-p\right]\;,$ (19)
@@
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 6: (a) Plot of cluster size distribution of perimeter bond clusters $n_{b}(p_{c})$ against the size of the perimeter bond cluster $b$ and cluster size distribution of nnn site clusters $n_{s}(p_{c})$ against nnn cluster size $s$ in a double logarithmic scale. (b) Plot of cluster size distribution of perimeter bond clusters $n_{b}(p)$ against perimeter cluster size $b$ in a double logarithmic scale for various values of $p$, indicated by different symbols, for a system size of $L=2048$. (c) Plot of $b_{\xi}$ and $s_{\xi}$ against $p-p_{c}$ in a double logarithmic scale. The value of the critical exponent $\sigma=2.52\pm 0.03$ was determined. (d) Plot of the collapse of ratio of cluster size distributions $n_{b}(p)/n_{b}(p_{c})$ against $(p-p_{c})b^{\sigma}$ for various values of $p$ at lattice size $L=2000$.
@@ 4.3 Cluster size distribution
$n_{b}(p_{c})\sim b^{-\tau}\;,$ (20)
as $\mathcal{N}[0]$ is a constant. In Fig. 6(a), the cluster size distribution $n_{b}(p_{c})$ of the perimeter bond cluster is plotted against the cluster size $b$ for the system of size $L=2000$ in the double logarithmic scale shown by the red squares. For comparison, the size distribution of the nnn site percolation clusters $n_{s}(p_{c})\sim s^{-\tau}$ is shown by blue triangles in Fig. 6(a). The slopes of the straight part of both the distributions are found to be $2.054\pm 0.002$. It is very close to the exponent $\tau=187/91$ as that of ordinary site percolation, shown by a black line as a guide to the eye. Note that the distributions $n_{b}(p_{c})$ and $n_{s}(p_{c})$ are two independent distributions. However, they only differ in the first bin and in the last few bins, though the total number of site and perimeter clusters are the same. In the first bin ($s\text{ or }b\in 1-10$), though there are all possible nnn site percolation clusters, the perimeter bond clusters appear in even numbers starting from $b=4$ only. The largest bin of $n_{b}(p_{c})$ is greater than that of $n_{s}(p_{c})$. This indicates that the perimeter of a site cluster can be more than the size of the site cluster.
Though the distribution is fairly power law for $p=p_{c}$, it deviates from the power law for $p<p_{c}$. The perimeter bond cluster size distributions for $p<p_{c}$ are plotted in Fig. 6(b) in double logarithmic scale. There is a strong deviation from the power law as $p$ is smaller than $p_{c}$. For a given $p$, if such a crossover occurs at a certain value of the perimeter bond cluster size $b_{\xi}(p)$, the cluster size distribution can be written in terms of $b_{\xi}$ as in sorensen
$n_{b}(p)=n_{b}(p_{c})F\left(\frac{b}{b_{\xi}}\right)\;,$ (21)
where $n_{b}(p_{c})\sim b^{-\tau}$. Here $b_{\xi}$ scales as
@@
where $\sigma$ is an exponent. We have measured $b_{\xi}$ from the intersection of the curve $\log(n_{b}(p))$ with the line $\log(n_{b}(p_{c}))$. To find the intersection point, we fit a higher order polynomial through $\log{(n_{b}(p))}$ around the point of intersection and find the solution with the line of $\log(n_{b}(p_{c}))$. Red crosses mark the $\log b_{\xi}$ values on the $x$-axis. In Fig. 6(c), $b_{\xi}$ is plotted against $|p-p_{c}|$ in double logarithmic scale and found that $b_{\xi}$ follows a power law with an exponent $1/\sigma=2.52\pm 0.03$ which is very close to the percolation value $1/\sigma=91/36\approx 2.53$. The corresponding quantity in nnn site percolation $s_{\xi}$ is also plotted for comparison. It seems the value of $\sigma$ is the same as that of the ordinary percolation.
@@
is a normalized scaling function of the scaled variable $z=(p-p_{c})b^{\sigma}$. The data for different $p$ for a given $L$ should collapse onto a single curve if $n_{b}(p)/n_{b}(p_{c})$ is plotted against $z$. In Fig. 6(d), the ratio $n_{b}(p)/n_{b}(p_{c})$ is plotted against the scaled variable $z=(p-p_{c})b^{\sigma}$ for several values of $p$ for the system of $L=2000$ taking $\tau=187/91$ and $\sigma=36/91$ as that of the ordinary site percolation . It can be seen that data for several $p$ have collapsed to a single curve with the $\tau$ and $\sigma$ of the ordinary site percolation. It can be noted that at $p=p_{c}$, one finds $z=0$ and that corresponds to $\mathcal{N}[0]/\mathcal{N}[0]=1$ on the $y$-axis. In terms of the parameter $\rho$, $n_{b}(\rho)$ is indifferent from $n_{b}(p)$. However, the normalized plot of $n_{b}(\rho)/n_{b}(\rho_{c})$ against the scaled variable $z=(\rho-\rho_{c})b^{\sigma}$ will have the same height but squeezed width.
Figure 7: Plot of mass of the spanning perimeter bond cluster $B_{\rm max}$ (on the dual lattice) against $L$ in black circles and plot of mass of the nnn spanning site cluster $S_{\rm max}$ (on the primal lattice) against $L$ in blue squares. The black and the blue lines of slope $91/48$ are guides to the eye.